What two factors of -375 that add up to 20
Answer by Edwin McCravy(20086)


It's not possible, and here is why:
(-375)(1) = -375, however, -375 + 1 = -374 not 20.
(-125)(3) = -375, however, -125 + 3 = -122 not 20.
(-75)(5) = -375, however, -75 + 5 = -70 not 20.
(-25)(15) = -375, however, -25 + 15 = -10 not 20.
(-15)(25) = -375, however, -15 + 25 = 10 not 20.
(-5)(75) = -375, however, -5 + 75 = 70 not 20.
(-3)(125) = -375, however, -3 + 125 = 122 not 20.
(-1)(375) = -375, however, -1 + 375 = 374 not 20.
(1)(-375) = -375, however, 1 + -375 = -374 not 20.
(3)(-125) = -375, however, 3 + -125 = -122 not 20.
(5)(-75) = -375, however, 5 + -75 = -70 not 20.
(15)(-25) = -375, however, 15 + -25 = -10 not 20.
(25)(-15) = -375, however, 25 + -15 = 10 not 20.
(75)(-5) = -375, however, 75 + -5 = 70 not 20.
(125)(-3) = -375, however, 125 + -3 = 122 not 20.
(375)(-1) = -375, however, 375 + -1 = 374 not 20.
Perhaps you meant to type 10 or 70 and typed 20 instead.
Edwin